OpenStudy (anonymous):

Part 1: Show all work in simplifying the quantity of three x squared plus three x minus thirty-six, divided by the quantity of x squared minus x minus six. Be sure to list restrictions. Part 2:Use complete sentences to explain the process.

OpenStudy (anonymous):

\[\frac{3x^2+3x-36}{x^2-x-6}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

factor out a 3 from the numerator first get \[\frac{3(x^2+x-12)}{x^2-x-6}\] then see if you can factor the top and bottom

OpenStudy (anonymous):

\[\frac{3(x+4)(x-3)}{(x+2)(x-3)}\] and then cancel

OpenStudy (anonymous):

clear now? restrictions are \(x\neq -2\) and \(x\neq 3\) because you cannot divide by zero

OpenStudy (anonymous):

Yess I see, thankss

OpenStudy (anonymous):

yw